What are three consecutive numbers that sums to 48? Which of these numbers are kewees 43,58,166,369,620,891, I need three numbers.
Answer by SwiftAlbatross(13)

Three numbers that sum to 48 could be represented by x , x + 1, and x + 2 so
(x) + (x + 1) + (x + 2) = 48
Combine like terms
x + x + x + 1 + 2 = 48
3x + 3 = 48
Subtract 3 from both sides
- 3 = - 3
3x = 45
Divide both sides by 3
/ 3 = / 3
x = 15
So the three numbers are 15, 16, and 17. You can check your work by summing them up.
15 + 16 + 17 = 48
Kewees are even numbers so 58, 166, and 620 are kewees.